# A Regret Perspective on Online Selective Generation

### v1

@@ Abstract
Large language generative models increasingly interact with humans, while their falsified responses raise concerns. To address this hallucination effect, selectively abstaining from answering, called selective generation, provides an effective way for generators to control the hallucination when it is unsure of their answers. However, as selective generators are interacting under non-stochastic environments and having partial feedback from users on selective generation (e.g., thumbs up or down on the selected answer), learning methods for selective generation under such practical setups are crucial but currently missing. To address these limitations, we propose an online learning algorithm for selective generation under partial feedback. In particular, as learning under partial feedback is well-studied by multi-armed bandit problems, we reduce selective generation to bandits and provide a novel conversion lemma from bandits back to selective generation to leverage any known bandit algorithms and theoretical properties. This mainly connects regret guarantees of bandits to false discovery rate (FDR) guarantees of selective generation for controlling hallucination. However, naively exploiting known bandit algorithms and their regret bounds suffers from slow convergence speed in practice due the nature of partial feedback. To overcome this, we exploit a unique structure of arms in selective generation for feedback unlocking, i.e., unlocking unknown feedback from observed feedback. We theoretically and empirically evaluate the efficacy of the proposed online selective generation algorithm under partial feedback over diverse data environment setups, resulting in controlling a desired FDR, while maintaining reasonable selection efficiency, i.e., the ratio of non-abstaining answers, compared to baselines.
@@ 1 Introduction
As large language generators openai2023gpt35turbo ; meta2024llama31-8binstruct surpass average human performance, their interaction with humans becomes increasingly prevalent. However, as they tend to generate incorrect information, or so-called hallucinated information, community’s concerns on their mis-alignment issues have significantly increased. One simple but effective approach to mitigating hallucination is selective generation, selectively abstaining from generation if the generator is unsure of its response geifman2017selective ; goren2024hierarchical ; mohri2023learning ; lee2024selective . By the abstention, the selective generator may control its high precision, or equivalently a low false discovery rate (FDR), i.e., whenever it says answers, they are mostly correct. Interestingly, few selective prediction and generation methods geifman2017selective ; goren2024hierarchical ; lee2024selective provide theoretical guarantees on the controllability of the FDR to ensure the trustworthiness of generators.
@@
However, conventional selective prediction methods geifman2017selective ; goren2024hierarchical ; lee2024selective are designed under limited stochastic assumptions, i.e., data are independently drawn from a fixed distribution, which undermines their applicability to real world applications in adversarial and distribution-shifted environments. Moreover, the methods require full feedback (e.g., a true answer is given) instead of partial feedback (e.g., the correctness of a generated answer), where partial feedback is more practical and easier to obtain (e.g., thumbs-up buttons in dialog systems).
@@
To mitigate these limitations, we propose a novel online selective generation algorithm to control a false discovery rate (FDR), while maximizing selection efficiency (i.e., the ratio of non-abstaining cases), with partial feedback. First, we address the online learning problem under partial feedback by leveraging profound achievements from online learning and multi-armed bandits. To this end, we reduce selective generation to bandits, exploit any regret minimization algorithms with regret bounds, and then translate the regret bounds back into FDR bounds for selective generation by introducing a novel Regret-to-FDR conversion lemma. This enables us to use any regret minimization algorithms for selective generation with FDR controllability guarantees.
@@
Second, we design a regret minimization algorithm tailored for selective generation to fully exploit feedback information. In particular, partial feedback is practical but the amount of new information is clearly insufficient compared to full feedback, suffering from sample inefficiency and slower convergence speed. To address this technical challenge, we exploit a unique structure of selective generators to unlock additional feedback from observed partial feedback, dubbed by partial feedback unlocking. By exploiting this, we extend the Exp3 algorithm auer2002nonstochastic for adversarial bandits for online Selective generation with partial feedback UnLocking (ExSUL). Moreover, we provide the $\mathcal{{O}}(\ell_{\text{max}}\sqrt{T\ln|\mathcal{{H}}|})$ regret bound of ExSUL, which is better than a regret minimizer with partial feedback, i.e., $\mathcal{{O}}(\ell_{\text{max}}\sqrt{T|\mathcal{{H}}|\ln|\mathcal{{H}}|})$, and compatible to a regret minimizer with full feedback, i.e., $\mathcal{{O}}(\ell_{\text{max}}\sqrt{T\ln|\mathcal{{H}}|})$, where $T$ is a time horizon and $\mathcal{{H}}$ is a set of hypotheses of selective generators. See Figure 1 for qualitative results of our method that controls an FDR in an interactive environment.
@@
We empirically evaluate the efficacy of the proposed online selective generation algorithm ExSUL. In particular, we consider two tasks (including question-answering over TriviaQA joshi2017triviaqa and Natural Question kwiatkowski2019natural and dialog conversations via two dialog agents), three learning environments (including stochastic, distribution-shifted, and interactive environments), and two language models (including GPT-3.5-turbo openai2023gpt35turbo and LLaMA3.1 meta2024llama31-8binstruct ) to show that our method (1) controls a desired FDR and (2) maintains reasonably low selection inefficiency, i.e., the ratio of abstention. Additionally, our learner’s FDR convergence speed is clearly better than a regret minimizer with partial feedback and compatible to a regret minimizer with full feedback (which forms a performance upper bound).
@@ 1.1 Related Work
Here, we introduce tightly related literature, including selective prediction, online learning, and bandit problems. See Appendix A for additional related work.
@@
Selective Prediction. Selective prediction is a method that abstains from making a prediction to control the FDR in a certified manner. geifman2017selective proposes selective classification that uses a threshold for a scoring function to abstain from uncertain predictions. goren2024hierarchical extends this to exploit hierarchical structure of labels for better efficiency. mohri2023learning extends this idea to generation tasks but mainly focuses on decontextualization task without the FDR guarantee. lee2024selective further generalizes selective prediction to generation tasks by introducing the concept of an entailment set, which addresses the challenge of open-ended question-answering. In particular, they propose a semi-supervised method to leverage unlabeled entailment data, providing a theoretical guarantee for controlling the FDR at a desired level. Previous work focuses on a batch learning setup under a stochastic assumption, which may be fragile in real-world applications. Contrast to this, we consider a non-stochastic setup, allowing distribution shift, to propose online selective generation methods.
Online Learning and Bandit Problems. Sequential prediction cesa2006prediction ; mohri2012foundations ; sankararaman2016semi ; foster2023foundations designs learners that adapt to sequentially arriving data. Online learning, e.g., exponential weighting littlestone1994weighted , mostly assumes adversarial environments where data can be arbitrarily, possibly adaptively, manipulated by adversaries. Under this environment, a learner receives full feedback on predictions. In contrast to this, bandit problems, including multi-armed bandits agrawal1995sample and adversarial bandits auer2002nonstochastic , assume to have sequentially-arriving partial feedback, i.e., a learner receives feedback on a chosen arm over time. To efficiently leverage partial feedback, structured bandits russo2013eluder exploit the functional structure of arms-to-loss functions and semi-bandits sankararaman2016semi allow to choose a set of arms with fixed size for better learning efficiency. We exploit online adversarial bandits to mitigate stochastic assumptions in learning traditional selective prediction and to consider practical learning under partial feedback. Moreover, we leverage a specific structure between arms and loss in selective generation for better learning efficiency.
2 Preliminary
We introduce preliminaries on language generation, selective prediction, and adversarial bandits. See Appendix B for online learning. To this end, let $\mathcal{{X}}$ be a set of inputs (e.g., examples or questions) and $\mathcal{{Y}}$ be a set of outputs (e.g., labels or answers).
2.1 Language Generation
We mainly consider language generators as our model to control hallucination. In particular, let $G:\mathcal{{X}}\to\mathcal{{Y}}$ be a language generator, where $\mathcal{{W}}$ is a set of tokens and $\mathcal{{X}}=\mathcal{{Y}}\coloneqq\cup_{i=0}^{\infty}\mathcal{{W}}^{i}$. Here, each $i$-th token $\hat{\mathbf{y}}_{i}$ of a generated answer $\hat{\mathbf{y}}\in\mathcal{{Y}}$ is decoded from an underlying probability distribution $p(\mathbf{y}\mid\mathbf{x})$, where $p$ is usually learned via language data. For the decoding strategy, we consider greedy decoding, i.e., $\hat{\mathbf{y}}_{i}=\arg\max_{w\in\mathcal{{W}}}p(w\mid\mathbf{x},\hat{\mathbf{y}}_{1:i-1})$, where $\mathbf{y}_{a:b}\coloneqq(\mathbf{y}_{a},\dots,\mathbf{y}_{b})$. Given a generated answer $\hat{\mathbf{y}}$, there are multiple ways to measure its likelihood of correctness $f:\mathcal{{X}}\times\mathcal{{Y}}\to\mathbb{R}$. A standard way $f_{\texttt{std}}$ considers a length-normalized token probability, i.e., $f_{\texttt{std}}(\mathbf{x},\hat{\mathbf{y}})\coloneqq(p(\hat{\mathbf{y}}_{1}\mid\mathbf{x})\prod_{i=2}^{|\hat{\mathbf{y}}|}p(\hat{\mathbf{y}}_{i}\mid\mathbf{x},\mathbf{\hat{y}}_{1:i-1}))^{\nicefrac{{1}}{{|\hat{\mathbf{y}}|}}}$. A better alternative is to consider consistency of $\hat{\mathbf{y}}$ to multiple generated answers with sampling manakul2023selfcheckgpt , i.e., $f_{\texttt{con}}(\mathbf{x},\hat{\mathbf{y}})=\hat{\mathbbm{E}}_{\mathbf{y}^{\prime}\sim G(x)}f_{E}(\mathbf{y}^{\prime},\hat{\mathbf{y}})$, denoted by a consistency score, where we use sampling for decoding $\mathbf{y}^{\prime}$, $f_{E}(\mathbf{y}^{\prime},\hat{\mathbf{y}})$ measures an entailment score of $\hat{\mathbf{y}}$ given $\mathbf{y}^{\prime}$ (i.e., whether $\mathbf{y}^{\prime}$ entails $\hat{\mathbf{y}}$) via an entailment model (e.g., GPT-3.5-turbo).
2.2 Selective Prediction
Selective prediction geifman2017selective and generation lee2024selective provide certified control over the risk of incorrect predictions by abstaining answers (saying IDK) when the prediction is uncertain. Given a predictor $\hat{\mathbf{y}}:\mathcal{{X}}\to\mathcal{{Y}}$, a selective predictor $\hat{{S}}:\mathcal{{X}}\to\mathcal{{Y}}\cup\{\texttt{IDK}\}$ abstains from returning $\hat{\mathbf{y}}(\mathbf{x})$ if a selection function $\hat{{s}}:\mathcal{{X}}\times\mathcal{{Y}}\to\{0,1\}$ deems the prediction uncertain, i.e., $\hat{{S}}(\mathbf{x})\coloneqq\begin{cases}\hat{\mathbf{y}}(\mathbf{x})&\text{if~{}}\hat{{s}}(\mathbf{x},\hat{\mathbf{y}}(\mathbf{x}))=1\\ \texttt{IDK}&\text{otherwise}\end{cases}.$ In learning, the selection function $\hat{{s}}$ is chosen to possibly satisfy a desired level of a false discovery rate, i.e., $\mathbbm{P}(\hat{{S}}(\mathbf{x})\neq_{E}\mathbf{y}\mid\hat{{S}}(\mathbf{x})\neq\texttt{IDK})$ over the i.i.d. samples of $(\mathbf{x},\mathbf{y})$. The equality relation $\neq_{E}$ is usually the standard exact matching in selective classification geifman2017selective , which is extended to capture semantic correctness via entailment in selective generation lee2024selective . In contrast to the standard stochastic setup as in the previous literature, we consider online learning under partial feedback.
2.3 Regret Minimization
In sequential prediction, the main goal is to find a learner whose performance is as good as the best learner in hindsight. As a goodness metric, we use widely accepted expected regret, simply called regret in this paper, leading to regret minimization. The goal of regret minimization is to find a distribution $p_{t}$ over hypotheses $\mathcal{{H}}$ to minimize regret, i.e., the gap between the learner’s expected cumulative loss over a possibly randomly chosen hypothesis $\tau_{t}\sim p_{t}$ and the best cumulative loss for any loss sequence $\ell_{t}$ over a time horizon $T$:
$\textbf{Reg}_{T}\coloneqq\sum_{t=1}^{T}\mathbbm{E}_{\tau_{t}\sim p_{t}}\ell_{t}(\tau_{t})-\min_{\tau\in\mathcal{{H}}}\sum_{t=1}^{T}\ell_{t}(\tau).$ (1)
Here, the best cumulative loss assumes to have access to loss for any $\tau$, i.e., full feedback, but the learner’s cumulative loss may have limited access to loss, e.g., partial feedback. Depending on this feedback setup, we consider two learning problems: online learning with full feedback and adversarial bandits with partial feedback. In the following section, we review adversarial bandits with one regret minimization algorithm, called Exp3. A well-known regret minimization algorithm for online learning with feedback is Exponential Weighting (EW) littlestone1994weighted , introduced in Appendix B.
2.4 Adversarial Bandits and Exp3 Algorithm
Traditional online learning typically assumes full feedback, i.e., loss for any hypothesis is computable. However, practical applications may only have partial feedback, i.e., loss for a chosen hypothesis is only given. To consider learning in partial feedback, we consider multi-armed bandit in an adversarial setting, called adversarial bandits. In adversarial bandits, we consider the same regret as in online learning, except that a learner only receives feedback of a chosen arm by itself, i.e., partial feedback, and an arbitrary loss sequence $\ell_{t}(\tau_{t})$ over a time horizon $T$ is fixed before learning by an oblivious adversary. In bandit context, we denote a hypothesis as an arm.
To minimize the regret $\textbf{Reg}_{T}$, Exponential-weight algorithm for Exploration and Exploitation (Exp3) auer2002nonstochastic is one standard algorithm, which leverages the EW algorithm for adversarial bandits (Algorithm 3). In particular, Exp3 exploits an unbiased estimator ~$\tilde{\ell}_{t}(\tau\mid\tau_{t})$ of loss for all arms $\tau\in\mathcal{{H}}$ constructed from loss $\ell_{t}(\tau_{t})$ of chosen arms $\tau_{t}$, i.e.,
~$\tilde{\ell}_{t}(\tau\mid\tau_{t})\coloneqq{\ell_{t}(\tau)}/{p_{t}(\tau)}\cdot\mathbbm{1}(\tau_{t}=\tau)\in[0,\infty).$ (2)
Note that ~$\tilde{\ell}_{t}(\tau\mid\tau_{t})$ is an unbiased estimator, i.e., $\mathbbm{E}_{\tau_{t}\sim p_{t}}\tilde{\ell}_{t}(\tau\mid\tau_{t})=\ell_{t}(\tau)$; thus, by using this in $\textbf{Reg}_{T}$, we have the following regret bound for Exp3. See Appendix G.2 for a proof.
Theorem 1.
auer2002nonstochastic ; foster2023foundations Let $\ell_{t}(\cdot)\in[0,\ell_{\text{max}}]$. For any $T\in\mathbb{N}$ and finite hypotheses $\mathcal{{H}}$, Algorithm 3 provides the following regret bound if $\eta=\sqrt{{2\ln|\mathcal{H}|}/{(\ell_{\max}^{2}T|\mathcal{H}|)}}$:
$\!\textbf{Reg}_{T}\!=\!\sum_{t=1}^{T}\mathbbm{E}_{\tau,\tau_{t}\sim p_{t}}\tilde{\ell}_{t}(\tau\mid\tau_{t})\!-\!\min_{\tau\in\mathcal{{H}}}\sum_{t=1}^{T}\ell_{t}(\tau)\!\leq\!\ell_{\max}\sqrt{2T|\mathcal{H}|\ln|\mathcal{H}|}\!=\!\mathcal{{O}}(\ell_{\text{max}}\sqrt{T|\mathcal{{H}}|\ln|\mathcal{{H}}|}).$ (3)
We will exploit Exp3 to extend this for learning selective generators with partial feedback.
@@ 3 Problem: Online Selective Generation with Partial Feedback
We consider online learning of a selective generator under partial feedback for language models. In particular, let $\mathcal{{W}}$ be a set of tokens and $\mathcal{{X}}=\mathcal{{Y}}\coloneqq\cup_{i=0}^{\infty}\mathcal{{W}}$ be a set of (token) sequences. Here, given a generator $G:\mathcal{{X}}\to\mathcal{{Y}}$, we consider a time-varying selective generator $\hat{{S}}_{t}:\mathcal{{X}}\to\mathcal{{Y}}\cup\{\texttt{IDK}\}$ that abstains from answering if a generated answer $G(\mathbf{x}_{t})$ at time $t$ is uncertain, i.e., $\hat{{S}}_{t}(\mathbf{x}_{t})\coloneqq\begin{cases}G(\mathbf{x}_{t})&\text{if~{}}\hat{{s}}(\mathbf{x}_{t},G(\mathbf{x}_{t}))=1\\ \texttt{IDK}&\text{otherwise}\end{cases},$ where $\hat{{s}}:\mathcal{{X}}\times\mathcal{{Y}}\to\{0,1\}$ is a selection function and IDK represents “I don’t know”. To learn the selective generator, we consider online learning under a non-stochastic assumption: each time step $t$ until a time horizon $T$, a learning algorithm (1) observes $\mathbf{x}_{t}\in\mathcal{{X}}$, (2) predicts $\hat{{S}}_{t}(\mathbf{x}_{t})$ where $\hat{{S}}_{t}$ is drawn from a learned distribution $p_{t}$ over selective generators, (3) observes partial feedback $b_{t}\coloneqq\mathbbm{1}(\hat{{S}}_{t}(\mathbf{x}_{t})\neq_{E}\mathbf{y}_{t})$ where $\mathbf{x}_{t}$ and $\mathbf{y}_{t}$ are drawn from a time-varying distribution, and (4) update the distribution $p_{t}$ over selective generators by using partial feedback. Here, $A\neq_{E}B$ means $A$ and $B$ are different in terms of a given correctness relation $E$, e.g., textual-entailment bowman2015large and $b_{t}=1$ if $\hat{{S}}_{t}(\mathbf{x}_{t})=\texttt{IDK}$ by the definition. Importantly, we consider partial feedback, i.e., we cannot directly observe $\mathbf{y}_{t}$ but we instead observe whether our prediction is correct or not, i.e., $b_{t}$. This is an ordinary setup in using language generators where a user provides thumb up or down feedback if a generated answer is correct or not, respectively.
Under learning with partial feedback, we aim to update a distribution $p_{t}$ over selective generators that controls an empirical false discovery rate (FDR), with respect to a correctness relation $E$ at a desired level $\alpha\in[0,1]$ up to a time horizon $T$, i.e.,
$\textbf{FDR}_{T}\coloneqq\frac{\sum_{t=1}^{T}\mathbbm{E}_{\hat{{S}}_{t}\sim p_{t}}\mathbbm{1}\left(\hat{{S}}_{t}(\mathbf{x}_{t})\neq\texttt{IDK}~{}\wedge~{}\hat{{S}}_{t}(\mathbf{x}_{t})\neq_{E}\mathbf{y}_{t}\right)}{\sum_{t=1}^{T}\mathbbm{E}_{\hat{{S}}_{t}\sim p_{t}}\mathbbm{1}\left(\hat{{S}}_{t}(\mathbf{x}_{t})\neq\texttt{IDK}\right)}\leq\alpha,$ (4)
where $\textbf{FDR}_{T}=\alpha$ if $\hat{{S}}_{t}(\mathbf{x}_{t})=\texttt{IDK}$ for all $t$ to penalize inefficient cases. Equivalently, we wish to find $\hat{{S}}_{1:T}$ that make the following FDR risk smaller than zero: $\sum_{t=1}^{T}\big{[}\mathbbm{1}(\hat{{S}}_{t}(\mathbf{x}_{t})\neq\texttt{IDK}\wedge\hat{{S}}_{t}(\mathbf{x}_{t})\neq_{E}\mathbf{y}_{t})-\alpha\mathbbm{1}(\hat{{S}}_{t}(\mathbf{x}_{t})\neq\texttt{IDK})\big{]}.$ Then, the goal is to learn a distribution $p_{t}$ over selective generators such that drawn selective generator $\hat{{S}}_{t}\sim p_{t}$ controls the FDR at a desired level $\alpha$ for any sequences, i.e.,
$\frac{1}{T}\sum_{t=1}^{T}\mathbbm{E}_{\hat{{S}}_{t}\sim p_{t}}\left[\mathbbm{1}\left(\hat{{S}}_{t}(\mathbf{x}_{t})\neq\texttt{IDK}\wedge\hat{{S}}_{t}(\mathbf{x}_{t})\neq_{E}\mathbf{y}_{t}\right)-\alpha\,\mathbbm{1}\left(\hat{{S}}_{t}(\mathbf{x}_{t})\neq\texttt{IDK}\right)\right]\leq\varepsilon(T)$ (5)
for any $\mathbf{x}_{t}$ and $\mathbf{y}_{t}$, where $\varepsilon(T)$ is some decreasing function in $T$. Additionally, we wish to minimize selection inefficiency $\textbf{Ineff}_{T}\coloneqq\frac{1}{T}\sum_{t=1}^{T}\mathbbm{E}_{\hat{{S}}_{t}\sim p_{t}}\mathbbm{1}(\hat{{S}}_{t}(\mathbf{x}_{t})=\texttt{IDK}).$
4 Method: Online Selective Generation with Partial Feedback Unlocking
We leverage a regret perspective from online learning and bandit problems in designing a learning algorithm for selective generation with full or partial feedback to control the FDR. In particular, we reduce selective generation to online learning with full feedback or bandit problems with partial feedback to use existing regret minimization algorithms. The main benefit of this reduction is that this enables us to leverage any regret minimization algorithms (e.g., EW or Exp3) and their regret bounds for selective generation algorithms with FDR bounds. To connect regret bounds and FDR bounds, we introduce a novel conversion lemma from the regret to the FDR in Section 4.4. However, under partial feedback setups, simply leveraging existing bandit algorithms may not lead to sample efficient algorithms without exploiting properties of selective generation.
In this section, we propose a novel online selective generation with partial feedback unlocking for sample efficiency. First, we introduce the reduction of the online selective generation to adversarial bandits in Section 4.1. In the following Section 4.2, we extend the Exp3 algorithm auer2002nonstochastic for adversarial bandits to learn selective generators by exploiting unique structure in selection functions by feedback unlocking, followed by its novel regret bound in Section 4.3. This regret bound leads to an FDR bound by using our conversion lemma in Section 4.4. Note that online selective generation under full feedback can be devised in a similar way. See Appendix C for this detail via EW.

### v2

# Online Selective Generation with Adversarial Bandit Feedback

@@ Abstract
Large language generative models increasingly interact with humans, while their falsified responses raise concerns. To mitigate this hallucination effect, selectively abstaining from answering, called selective generation, provides an effective way for generators to control the hallucination when uncertain about their answers. However, as selective generators interact under adversarial environments and receive partial feedback from users on selected generation (e.g., thumbs up or down on the selected answer), learning methods for selective generation under such practical setups are crucial but currently missing. To address this limitation, we propose an online learning algorithm for selective generation with partial feedback under an adaptive adversary. In particular, we re-purpose an adversarial bandit algorithm to design an online selective generation method with controllable false discovery rates (FDR), which measures the rate of hallucination. The key building blocks include a novel conversion lemma from regret of any bandit algorithm to the FDR, and the exploitation of a unique structure of selective generation to reuse partial feedback, which we call feedback unlocking. We empirically evaluate the efficacy of the proposed online selective generation algorithm with partial feedback over diverse learning environments, demonstrating its ability to control the FDR, while maintaining reasonable selection efficiency, i.e., the ratio of non-abstaining answers, compared to baselines.
@@ 1 Introduction
As large language generators (OpenAI, 2023; Meta AI, 2024) surpass average human performance, their interaction with humans becomes increasingly prevalent. However, their tendency to generate incorrect information, or so-called hallucination, has significantly raised community concerns on their mis-alignment issues. One simple but effective approach to mitigating hallucination is selective generation, selectively abstaining from generation if the generator is unsure of its response (Geifman & El-Yaniv, 2017; Goren et al., 2024; Mohri et al., 2023; Lee et al., 2024). By the abstention, the selective generator may control its high precision, or equivalently a low false discovery rate (FDR), i.e., whenever it says answers, they are mostly correct. Interestingly, few selective prediction and generation methods (Geifman & El-Yaniv, 2017; Goren et al., 2024; Lee et al., 2024) provide theoretical guarantees on the controllability of the FDR to ensure the trustworthiness of generators.
@@
However, conventional selective prediction methods (Geifman & El-Yaniv, 2017; Goren et al., 2024; Lee et al., 2024) are designed under limited stochastic assumptions, i.e., data are independently drawn from a fixed distribution, which undermines their applicability to real world applications in adversarial and distribution-shifted environments. Moreover, the methods require full feedback (e.g., a true answer is given) instead of partial feedback (e.g., the correctness of a selectively generated answer), where partial feedback is more practical and easier to obtain (e.g., thumbs-up buttons in dialog systems). To mitigate these limitations, we propose a novel online selective generation algorithm to control a false discovery rate (FDR), while maximizing selection efficiency (i.e., the ratio of non-abstaining cases), with partial feedback.
@@
First, we address online selective generation under partial feedback by leveraging profound achievements from online learning and multi-armed bandits. To this end, we reduce online selective generation to bandits and exploit any regret minimization algorithms, by introducing a novel Regret-to-FDR conversion lemma.
@@
Second, we design an algorithm tailored for selective generation which fully exploits feedback information. Although partial feedback is practical, it provides the insufficient amount of information compared to full feedback, leading to sample inefficiency and slower convergence speed. To address this challenge, we exploit the unique structure of selective generation to unlock additional information from partial feedback, a technique we call partial feedback unlocking. Leveraging this idea, we extend the Exp3-IX algorithm (Neu, 2015) for adversarial bandits to online Selective generation with partial feedback UnLocking (ExSUL). Moreover, we provide the $\mathcal{{O}}(\sqrt{T\ln|\mathcal{{H}}|})$ expected regret bound of ExSUL—which affects the convergence of the FDR guarantee—better than a learner with partial feedback, i.e., $\mathcal{{O}}(\sqrt{T|\mathcal{{H}}|\ln|\mathcal{{H}}|})$, and compatible to that with full feedback, i.e., $\mathcal{{O}}(\sqrt{T\ln|\mathcal{{H}}|})$, where $T$ is a time horizon and $\mathcal{{H}}$ is a set of hypotheses of selective generators.
@@
Finally, we empirically evaluate the efficacy of the proposed online selective generation algorithm ExSUL. In particular, we consider two tasks (question-answering over TriviaQA (Joshi et al., 2017) and Natural Question (Kwiatkowski et al., 2019) and dialog conversations via two dialog agents), three learning environments (stochastic, distribution-shifted, and interactive), and two language models (GPT-3.5-turbo (OpenAI, 2023) and LLaMA3.1 (Meta AI, 2024)), demonstrating that our method (1) controls the desired FDR and (2) maintains reasonably low selection inefficiency, i.e., the ratio of abstention. See Figure 1 for qualitative results that control an FDR in an interactive environment.
@@ 1.1 Related Work
Here, we introduce closely related literature, including selective prediction, online learning, and bandit problems. See Appendix C for additional related work.
@@
Selective Prediction. Selective prediction is a method that abstains from making a prediction to control the FDR in a certified manner. (Geifman & El-Yaniv, 2017) proposes selective classification, which abstains from uncertain predictions by learning a threshold for a scoring function. (Goren et al., 2024) extends this framework by leveraging the hierarchical structure of classification labels to improve efficiency. (Mohri et al., 2023) adapts this idea to generation tasks, though the focus is primarily on the decontextualization without providing an FDR guarantee. (Lee et al., 2024) further generalizes selective prediction to generation tasks by introducing the concept of an entailment set, which addresses the challenge of open-ended question-answering. In particular, they propose a semi-supervised method to leverage unlabeled entailment data, providing an FDR guarantee. However, prior work mainly focuses on a batch learning under a stochastic assumption, i.e., i.i.d. assumption on data distributions, which may be fragile in real-world applications. In contrast, we consider a non-stochastic setup that allows for distribution shift and propose online selective generation methods.
Online Learning and Bandit Problems. Sequential prediction (Cesa-Bianchi & Lugosi, 2006; Mohri et al., 2012; Sankararaman, 2016; Foster & Rakhlin, 2023) designs learners that adapt to sequentially arriving data. Unlike stochastic learning, online learning methods (e.g., exponential weighting (Littlestone & Warmuth, 1994)) make minimal distributional assumptions and analyze worst-case, possibly adaptive sequences, with guarantees in terms of regret minimization. In the full-information setting, the learner observes the loss of every action each round.
Furthermore, bandit problems, such as multi-armed bandits (Agrawal, 1995) and adversarial bandits (Auer et al., 2002; Bubeck et al., 2012), address a partial-feedback setting, where the learner only observes feedback of the chosen action (arm) each round. A well-known algorithm is Exp3 (Auer et al., 2002), which leverages exponential weighting (Littlestone & Warmuth, 1994) with an importance-weighted unbiased loss estimator. However, Exp3 suffers from a dishearteningly large variance of the regret, leading to looser concentration. To obtain high-probability bounds, variants introduce exploration explicitly or implicitly, i.e., Exp3.P (Auer et al., 2002; Bubeck et al., 2012) explicitly mixes the uniform distribution, and Exp3-IX (Neu, 2015) implements implicit exploration by adding an exploration parameter to the loss estimator. To efficiently leverage partial feedback, structured bandits (Russo & Van Roy, 2013) exploit the functional structure of arms-to-loss functions and semi-bandits (Sankararaman, 2016) allows to choose a set of arms with fixed size for better learning efficiency.
In particular, we exploit adversarial bandits to mitigate stochastic assumptions in traditional selective prediction and to consider learning under partial feedback. Moreover, we leverage a unique structure between arms and loss in selective generation to achieve better learning efficiency.